Vestport, Inc.has 20 million common shares outstanding priced at $25.00 each. Next year’s dividend on these shares is expected to be $1.50, and the dividend is expected to grow at 5% per year forever. Vestport has 150,000 bonds outstanding, each with a $1,000 par value, 10% annual coupon rate (payable semiannually), 10 years to maturity, and priced at $1,050 . Vestport's marginal corporate income tax rate is 21%. Compute the WACC for the Vestport, Inc. PLEASE SELECT ALL OF THE FOLLOWING STATEMENTS THAT ARE TRUE.

Group of answer choices

The before-tax cost of debt is 9.22% annually

The after-tax cost of equity is 11%

The weighted average cost of capital is 10.11%

The market value proportion of debt financing is less than 25%
Answer #1

ALL THE STATEMENTS ARE TRUE

Before-tax cost of debt=RATE(10*2,10%*1000/2,-1050,1000)*2=9.224%

After-tax cost of equity=D1/P0+g=1.5/25+5%=11.000%

WACC=(20*25*10^6*11%+150000*1050*9.224%*(1-21%))/(20*25*10^6+150000*1050)=10.111%

Market value proportion of debt financing=150000*1050/(20*25*10^6+150000*1050)=23.954%

Hence, all the statements are true